NaCl + H2SO4 —> NaHSO4 + HCl If a reaction vessel contains 10.0 g of sodium chloride and 12.0 g of sulfuric acid, what is the limiting reactant? 


1
Expert's answer
2021-05-31T02:08:47-0400

NaCl + H2SO4 —> NaHSO4 + HCl

n = m / M

M (NaCl) = 58.4 g/mol

M (H2SO4) = 98.1 g/mol

n (NaCl) = 10 / 58.4 = 0.17 mol

n (H2SO4) = 12 / 98.1 = 0.12 mol

According to the equation, n(NaCl) = n(H2SO4). From the given reactant amounts, H2SO4 is the limiting reactant.